OpenStudy (anonymous):

Solve using the Quadratic Formula. 4x² + 7x + 1 = 0

OpenStudy (anonymous):

Please help!! this test is determing if i pass or not

sam (.sam.):

\[4x^2 + 7x + 1 = 0\] Written in \[ax^2+bx+c=0\] The quadratic formula is \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] substitute in the values a is 4 b is 7 c is 1

OpenStudy (anonymous):

A moment.

OpenStudy (anonymous):

Im still confused!

OpenStudy (whpalmer4):

What exactly is your point of confusion? Do you not understand how to evaluate the formula, or how it is derived, or ???

OpenStudy (luigi0210):

laziness

OpenStudy (anonymous):

am i sopose to plug in those values and no im not lazy when i ask question i need help not to answer them for me

OpenStudy (anonymous):

Substitute the respective formulas in the equation,it's the simplest.

OpenStudy (whpalmer4):

.Sam. gave you the formula, and told you to substitute the numbers for each letter (even identifying the values of each letter).

OpenStudy (whpalmer4):

Your equation is \(4x^2+7x+1=0\) so \(a = 4, b = 7, c = 1\) and the formula is \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] Wherever you see \(a\) substitute 4; \(b\) substitute 7; \(c\) substitute 1

OpenStudy (anonymous):

ok and i got -7+ or - \[\sqrt{33}\] over 8

OpenStudy (whpalmer4):

Very good...

OpenStudy (anonymous):

thank you..see im not lazy i just need some help i didnt ask for the answer i asked you to help me

OpenStudy (anonymous):

x = 1/8 (-7-sqrt(33))

OpenStudy (whpalmer4):

I didn't accuse you of laziness, I just asked what was confusing you... @Odero there are two solutions because the polynomial has an x^2 term. the other solution is x = 1/8(-7+sqrt(33))

OpenStudy (anonymous):

Ooops,thanks...